OpenStudy (anonymous):

If j, k, and n are consecutive integers such that 0 less than j less than k less than n and the units (ones) digit of the product j times n is 9, what is the units digit of k

OpenStudy (anonymous):

\(j,k,n\) are consecutive integers, so if \(j\) is the smallest, then \(k=j+1\) and \(n=j+2\). The product \(jn\) has a 9 in the ones digit, which means the ones digits of these two numbers are factors of 9, either 3 and 3, or 9 and 1. Suppose the ones digit of both \(n\) and \(j\) is 3. To simplify, suppose \(j=3\). But then \(n=j+2=5\), which contradicts our assumption. This means the ones digit of \(j\) is either 1 or 9, and for \(n\) it's either 9 or 1. Clearly, we have the first case. For instance, if \(j=9\), then \(n=j+2=11\) (note the ones digits). This means \(k=j+1=10\), and so the ones digit of \(k\) is 0.